OpenStudy (anonymous):

Solve the equation by a method of your choice: 6x2 + 5x + 1 = 0 If there are two solutions, separate the solutions by a comma.

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (noelgreco):

Use grouping of the linear term

OpenStudy (mrnood):

(ax+b)(cx+d) you can see that b and c MUST be 1 to multiply out to 1 so (ax + 1)(bx+1) find a and b so ab=6 and a+b = 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what do i do 6 and 5

OpenStudy (anonymous):

@BunnyBree

OpenStudy (mrnood):

I have shown you haow to factoise the equagtion, and shown that the constant term must be 1 in each bracket so (ax + 1)(bx+1)=6x2 + 5x + 1 so a*b =6 a+b = 5 find the values of a and b (you might just need to think about it for a minute...)

OpenStudy (anonymous):

ok um i dont know

OpenStudy (mrnood):

OK think of 2 numbers that multiply toghether to make 6 (there are only 2 possible pairs - one of them is 1 and 6) think of the other one. Do those two numbers add up to 5?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@AriPotta

OpenStudy (mrnood):

what are the other 2 numbers that multiply to make 6 (other than 1 *6)?